# Self-compensating Light Calorimetry with Liquid Argon Time Projection Chamber for GeV Neutrino Physics

### v1

@@ IV Self-compensating light calorimetry
The features of the LArTPC light calorimetry seen in Fig. 3 can be understood in analogous to the compensating calorimeter concept in HEP accelerator experiments, which is an important technique because of the fundamental problems of hadron calorimetry [1, 42]. The development of hadronic cascades results in both the electromagnetic (EM) component and the hadronic component. The EM component ($e$) comes from the $\pi^{0}$ and $\eta$ particles generated in the hadronic interactions, which decay into $\gamma$-rays that develop into EM showers. The hadronic component ($h$) consists of everything else such as protons, neutrons, and charged pions. The calorimeter response ($R_{\mathrm{cal}}$), defined as the ratio of the visible energy to the available energy, is very different between $e$ and $h$. As described in Sec. III.2, missing energy due to nuclear breakup can be up to 30%–50% for hadrons with large event-by-event fluctuation, while for EM particles the response is typically higher with smaller fluctuations. A calorimeter is characterized by the $e/h$ ratio, defined as the ratio between $R_{\mathrm{cal}}^{e}$ and $R_{\mathrm{cal}}^{h}$. Calorimeters with $e/h=1$ are called compensating calorimeters. For non-compensating calorimeters, the response to the total energy depends on the fraction of visible energy in the EM component ($f_{em}$), which has a large event-by-event fluctuation and often dominates the calorimeter energy resolution.
@@
Compensation between the $e$ and $h$ responses is typically realized in sampling calorimeters using a combination of three techniques. First, the $e$ response can be reduced by using high-Z absorber materials. Second, the neutron response, which is a significant component in $h$, can be boosted by using hydrogenous active materials. Finally, the sampling fraction can be tuned to match $e$ and $h$ responses so that $e/h=1$. Examples of high-performance compensating colorimeters include the ZEUS calorimeter at HERA using uranium absorbers and plastic scintillators [43], and the Spaghetti Calorimeter (SPACAL) at CERN using lead absorbers and scintillation fibers [44].
@@
Figure 4: The calorimeter responses ($R_{\mathrm{cal}}$, the ratio of the visible energy to the available energy) in LArTPC for the EM component ($e$, blue) and the hadronic component ($h$, red). The top panel shows the charge calorimetry and the bottom panel shows the light calorimetry. $10^{5}$ $\nu_{e}$ CC events are simulated with neutrino energy of 0.5–5 GeV. The area of each distribution is normalized to 1.
@@
In a LArTPC neutrino detector, the $\nu_{e}$ interaction (Eq. 11) guarantees there are both EM (electrons and possibly $\pi^{0}$s) and hadronic components in the calorimeter. The fraction of EM component ($f_{\mathrm{em}}$) varies event-by-event and is energy-dependent, which is modeled by the neutrino event generator such as GENIE. Similar to the problems of hadron calorimeters mentioned previously, if the LArTPC’s calorimeter response ($R_{\mathrm{cal}}$) between $e$ and $h$ is very different, the fluctuation of $f_{\mathrm{em}}$ will largely impact the calorimeter energy resolution and the shape of the energy response. The calorimeter responses in LArTPC are studied with the detector simulation described in Sec. III.2. For each $\nu_{e}$ event, the electron and $\pi^{0}$s, if any, are grouped into $e$, and the protons, neutrons, pions, and other final-state particles are grouped into $h$. $R_{\mathrm{cal}}$ in each component is calculated and plotted in Fig. 4. The top panel shows the $e$ (blue) and $h$ (red) responses in the charge calorimetry and the bottom panel shows the light calorimetry. The event-by-event fluctuation in the $e$ responses is much smaller than the $h$ responses because of the much smaller contribution of missing energy in $e$. The most interesting feature, however, is that the peak of $e/h$ in the light calorimetry is close to 1, which makes LArTPC a self-compensating light calorimeter.
@@
Figure 5: A scatter plot of 1000 simulated 3 GeV $\nu_{e}$-Ar CC events in a LArTPC. Activities in each event are grouped into the EM ($e$, solid circles) and hadronic ($h$, crosses) components. The y-axis shows the average recombination factor for charge ($\overline{\beta R_{c}}$, top) or light ($\overline{R_{L}}$, bottom) in $e$ (blue solid circles) or $h$ (red crosses). The x-axis shows the average ratio of the deposited energy to the total available energy in $e$ or $h$. The product of x and y is the calorimeter response shown in Fig. 4, where the $xy=0.42$ curve corresponds to the peak position of the $e$ or $h$ distribution in the light calorimeter response (bottom panel of Fig. 4).
@@
Figure 6: (Top) The light recombination factor ($R_{L}$) as a function of the drift electric field. The three curves correspond to particles with $dE/dx$ equaling to 1, 2, and 3 times that of a minimum ionizing particle (2.1 MeV/cm), respectively. (Bottom) The calorimeter response ratio between $e$ and $h$ ($e/h$) as a function of the drift electric field.
@@
The compensation performance of the LArTPC light calorimetry is dependent on the size of the drift electric field. The top panel of Fig. 6 shows the light recombination factor ($R_{L}$) as a function of the drift electric field derived from Eq. 6 and 7. The different colored curves show the $R_{L}$ for particles with $dE/dx$ equaling to 1, 2, and 3 times that of a MIP ($dE/dx=2.1$ MeV/cm). At a fixed drift E-field, the more separated are $R_{L}$s, the more compensation there is between the $e$ and $h$ responses. The $e/h$ is calculated at each E-field, and the optimal compensation of $e/h$ close to 1 can be achieved between 0.3–0.5 kV/cm. The compensation is slightly worse at lower and higher E-fields, but $e/h$ between 1–1.05 can still be achieved across a broad range of E-fields between 0.2–1.8 kV/cm.
@@
Figure 7: The calorimeter response of the hadronic component $h$ in the light calorimetry at different incident neutrino energies.
@@
The compensation performance in the light calorimetry is also energy-dependent. To study the energy dependence, the events in Fig. 4 are divided into different incident neutrino energy bins, and the calorimeter response of the hadronic component ($R_{\mathrm{cal}}^{h}$) in the light calorimetry is plotted in Fig. 7. For neutrino energies greater than 2 GeV, the $R_{\mathrm{cal}}^{h}$ distributions are similar, with smaller fluctuations for higher energy neutrinos. The $R_{\mathrm{cal}}^{h}$ distribution, however, is much wider for the 1 GeV or lower-energy incident neutrinos, which leads to a worse reconstructed energy resolution for lower-energy neutrinos (cf. Fig. 9) using light calorimetry.
@@
Figure 8: The calorimeter response ($R_{\mathrm{cal}}$) for protons, neutrons, and charged pions in the light calorimetry. The top panel shows a scatter plot of the available energy of each particle versus its $R_{\mathrm{cal}}$. The bottom panel shows the projected distribution of $R_{\mathrm{cal}}$ of different particles for all energies.
@@
This energy dependence is further studied in Fig. 8 by separately plotting the $R_{\mathrm{cal}}$ for protons, neutrons, and charged pions in the light calorimetry. The top panel shows a scatter plot of the available energy of each particle (i.e. the kinetic energy of protons and neutrons, and the total energy of the charged pions) versus its $R_{\mathrm{cal}}$. The bottom panel shows the projected distribution of $R_{\mathrm{cal}}$ of different particles for all energies. We can observe that at low energy below $\sim$1 GeV, the behavior is quite different between $p$, $n$, and $\pi^{\pm}$. This is because a large contribution of $R_{\mathrm{cal}}$ comes from the primary particle itself before it possibly starts a hadronic shower. Proton has a higher $R_{\mathrm{cal}}$ because it starts as a charged particle without missing energy, and receives extra recombination light. Neutron has a lower $R_{\mathrm{cal}}$ because it has a high missing energy from the beginning. Charge pion has a similar $dE/dx$ as a MIP particle before it starts the hardronic shower, so its $R_{\mathrm{cal}}$ is similar to electrons, with additional missing energy from particle decays. Therefore, the total $R_{\mathrm{cal}}$ of the hadronic component becomes dependent on the distribution of final state particles, leading to a wider distribution for lower energy neutrinos as seen in Fig. 7. However, at higher energy above $\sim$2 GeV, the $p$, $n$, and $\pi^{\pm}$ behave very similarly because they all start hadronic cascades and the $R_{\mathrm{cal}}$ is dominated by the secondary shower particles, which are similar for different primary particles. Thus, the compensation performance is better for higher-energy neutrinos and less sensitive to the modeling of final-state particles after neutrino interactions.
V Energy resolution comparision
@@
Based on the discussions in Sec. IV and Appendix C, in this section, we describe four energy reconstruction methods in LArTPC, one for light calorimetry and three for charge calorimetry, and compare their performance. The light calorimetry utilizes the self-compensating calorimeter concept first described in this paper. The charge calorimetry methods are similar in concept to previous works from LArTPC experiments and are included here for comparison.
@@
    L1: Simple light calorimetry by summing up $L$ from all particle activities and scaling to the incident neutrino energy:
@@
    Q1: Simple charge calorimetry by summing up $Q$ from all particle activities and scaling to the incident neutrino energy:
@@
    Q2: Assuming $e$ and $h$ charge activities can be grouped using pattern recognition algorithms on the 3D image (cf. Appendix C), scaling the $Q$ from $e$ and $h$ separately before summing up to the incident neutrino energy:
@@
We also studied two additional reconstruction methods but they are not reported in detail here. The first method is the $Q+L$ calorimetry based on Eq. 8. While it removes the recombination effect and recovers $E_{\mathrm{dep}}$, it does not account for the missing energy contribution, therefore, its performance is only slightly better than Q1 but worse than Q2 and L1. The second method assumes activities from individual particles can be grouped and their energies can be reconstructed separately. This is mostly useful for muons and charged pions since their missing energies from decays or captures can be corrected event-by-event, using methods such as a range-based energy reconstruction. The missing energy fluctuation from nuclear breakup still remains, e.g. from neutrons, making its performance only slightly better than Q3.
@@
Figure 9: (top) The distribution of the reconstructed neutrino energy from $10^{4}$ simulated 1 GeV $\nu_{e}$-Ar CC interaction events. The four colored histograms correspond to the four reconstruction methods Q1–3 and L1, respectively. (bottom) The relative energy resolution, $\sigma/\bar{E}$, as a function of the incident neutrino energy for the four reconstruction methods Q1–3 and L1.
@@
The top panel of Fig. 9 shows the distribution of the reconstructed neutrino energy from $10^{4}$ simulated 3 GeV $\nu_{e}$-Ar CC interactions using the Q1–3 and L1 reconstruction methods. Q1 (black) has a long tail extending to low energies, which comes from the missing energy contribution and the different calorimeter response between $e$ and $h$. L1 (red) removes the low-energy tail through the self-compensation mechanism, resulting in a more symmetric shape and much better resolution. Q2 (green) has a similar performance as L1, sharper near the peak but with a longer tail in the high energy. Q3 (blue) has the best overall performance. Both Q2 and Q3 utilize LArTPC’s charge imaging capability to separate $e$ and $h$ activities to improve the reconstructed energy resolution.
@@
The bottom panel of Fig. 9 shows the relative energy resolution, $\sigma/\bar{E}$, as a function of the incident neutrino energy. For each fixed neutrino energy, the $\sigma$ is calculated from the root mean square (RMS) of the distribution from each reconstruction method, and the mean energy of the distribution is used as $\bar{E}$. The RMS is larger than the width of the peak for a tailed distribution such as Q1 and Q2. This plot shows that Q3 achieves the best energy resolution at $\sim$6at 1 GeV and Q1 has the worst resolution of $\sim$14at 1 GeV. The L1 method has an 8.4% resolution at 1 GeV. The Q2 method performs similarly to L1, although the longer tail at high energy makes its RMS worse than L1. In practice, the 2D energy response matrices of reconstructed neutrino energy vs. true neutrino energy, also called the detector smearing matrices, are used in physics analysis, and they are shown in Fig. 11 in Appendix D. The energy resolution and response matrices from this study can be readily compared with other literature or experimental works [45, 38, 22, 46, 47], and can be used in the phenomenology studies such as those in Ref. [45, 48]. One should keep in mind the experimental conditions assumed in different studies.
@@
It’s worth noting that although Q3 has the best overall performance, the light calorimetry L1 is a much simpler reconstruction method without needing complex pattern recognition algorithms to separate $e$ versus $h$, or track-like versus dot-like activities. For higher energy neutrinos above 3 GeV, the performance between L1 and the much more sophisticated Q3 methods nearly equals owing to the self-compensation mechanism in the light calorimetry. Therefore, the light calorimetry in LArTPC provides an independent and comparable energy estimation, complementing the well-established charge calorimetry. The synergy between the light and charge calorimetry offers a novel approach to evaluating and mitigating the systematic uncertainties in energy measurements with LArTPCs.

### v2

@@ IV Self-compensating light calorimetry
The features of the LArTPC light calorimetry seen in Fig. 3 can be understood in analogous to the compensating calorimeter concept in HEP accelerator experiments, which is an important technique because of the fundamental problems of hadron calorimetry [1, 44]. The development of hadronic cascades results in both the electromagnetic (EM) component and the hadronic component. The EM component ($e$) comes from the $\pi^{0}$ and $\eta$ particles generated in the hadronic interactions, which decay into $\gamma$ rays that develop into EM showers. The hadronic component ($h$) consists of everything else such as protons, neutrons, and charged pions. The calorimeter response ($R_{\mathrm{cal}}$), defined as the ratio of the visible energy to the available energy, is very different between $e$ and $h$. As described in Sec. III.2, missing energy due to nuclear breakup can be up to 30%–50% for hadrons with large event-by-event fluctuation, while for EM particles the response is typically higher with smaller fluctuations. A calorimeter is characterized by the $e/h$ ratio, defined as the ratio between $R_{\mathrm{cal}}^{e}$ and $R_{\mathrm{cal}}^{h}$. Calorimeters with $e/h=1$ are called ideal compensating calorimeters. The benefit of compensating calorimeters comes from the principle of calorimeter energy reconstruction:
$E_{\mathrm{rec}}=\frac{f_{em}\cdot E_{\mathrm{vis}}}{R_{\mathrm{cal}}^{e}}+\frac{(1-f_{em})\cdot E_{\mathrm{vis}}}{R_{\mathrm{cal}}^{h}},$ (12)
where $f_{em}$ is the fraction of the visible energy in the EM component and $E_{\mathrm{rec}}$ is the reconstructed energy of the incident particle. Since $f_{em}$ has a large event-by-event fluctuation and is difficult to measure, it often dominates the calorimeter energy resolution. For ideal compensating calorimeters with $e/h=1$, however, Eq. 12 can be simplified to $E_{\mathrm{rec}}=E_{\mathrm{vis}}/R_{\mathrm{cal}}$, which does not suffer from $f_{em}$ fluctuations.
@@
Compensation between the $e$ and $h$ responses is typically realized in sampling calorimeters using a combination of three techniques. First, the $e$ response can be reduced by using high-Z absorber materials. Second, the neutron response, which is a significant component in $h$, can be boosted by using hydrogenous active materials. Finally, the sampling fraction can be tuned to match $e$ and $h$ responses so that $e/h$ is close to 1. Examples of high-performance compensating colorimeters include the ZEUS calorimeter at HERA using uranium absorbers and plastic scintillators [45], and the Spaghetti Calorimeter (SPACAL) at CERN using lead absorbers and scintillation fibers [46].
@@
Figure 5: Calorimeter responses ($R_{\mathrm{cal}}$, the ratio of the visible energy to the available energy) in LArTPC at the drift E-field of 0.5 kV/cm for the EM component ($e$, blue) and the hadronic component ($h$, red). The top panel shows the charge calorimetry and the bottom panel shows the light calorimetry. $10^{5}$ $\nu_{e}$ CC events are simulated with a uniform neutrino energy distribution between 0.5–5 GeV. The area of each distribution is normalized to 1.
@@
In a LArTPC neutrino detector, the $\nu_{e}$-Ar CC interaction (Eq. 11) guarantees there are both EM (electrons and possibly $\pi^{0}$s) and hadronic components in the calorimeter. The fraction of the visible energy in the EM component ($f_{\mathrm{em}}$) varies event by event and is energy-dependent, which is modeled by the neutrino event generator such as GENIE. Similar to the problems of hadron calorimeters mentioned previously, if the LArTPC’s calorimeter response ($R_{\mathrm{cal}}$) between $e$ and $h$ is very different, the fluctuation of $f_{\mathrm{em}}$ will largely impact the calorimeter energy resolution and the shape of the energy response. The calorimeter responses in LArTPC are studied with the detector simulation described in Sec. III.2. For each $\nu_{e}$-Ar CC event, the electron and $\pi^{0}$s, if any, are grouped into $e$, and the protons, neutrons, pions, and other final-state particles are grouped into $h$. $R_{\mathrm{cal}}$ in each component is calculated and plotted in Fig. 5. The top panel shows the $e$ (blue) and $h$ (red) responses in the charge calorimetry and the bottom panel shows the light calorimetry. The event-by-event fluctuation in the $e$ responses is much smaller than the $h$ responses because of the much smaller contribution of missing energy in $e$. The most interesting feature, however, is that the peak of $e/h$ in the light calorimetry is close to 1, which makes it a compensating calorimeter. Since no extra material needs to be added to achieve compensation between the $e$ and $h$ responses, we call LArTPC a self-compensating light calorimeter.
@@
Figure 6: Scatter plot of 1000 simulated 3 GeV $\nu_{e}$-Ar CC events in a LArTPC at the drift E-field of 0.5 kV/cm. Activities in each event are grouped into the EM ($e$, solid circles) and hadronic ($h$, crosses) components. The y-axis shows the average recombination factor for charge ($\overline{\beta R_{c}}$, top) or light ($\overline{R_{L}}$, bottom) in $e$ (blue solid circles) or $h$ (red crosses). The x-axis shows the average ratio of the deposited energy to the total available energy in $e$ or $h$. The product of x and y is the calorimeter response shown in Fig. 5, where the $xy=0.42$ curve corresponds to the peak position of the $e$ or $h$ distribution in the light calorimeter response (bottom panel of Fig. 5).
@@
Figure 7: (Top) Light recombination factor ($R_{L}$) as a function of the drift electric field. The three curves correspond to particles with $dE/dx$ equal to one, two, and three times that of a minimum ionizing particle (2.1 MeV/cm), respectively. (Bottom) Simulated calorimeter response ratio between $e$ and $h$ ($e/h$) as a function of the drift electric field. For each drift electric field from 0.1–2 kV/cm (with 0.1 kV/cm increments), $10^{4}$ $\nu_{e}$-Ar CC events are simulated with a uniform neutrino energy distribution from 0.5–5 GeV.
@@
The compensation performance of the LArTPC light calorimetry is dependent on the size of the drift electric field. The top panel of Fig. 7 shows the light recombination factor ($R_{L}$) as a function of the drift electric field derived from Eqs. 6 and 7. The different colored curves show the $R_{L}$ for particles with $dE/dx$ equal to one, two, and three times that of a MIP ($dE/dx=2.1$ MeV/cm). At a fixed drift electric field, the more separated are the $R_{L}$s, the more compensation there is between the $e$ and $h$ responses. To quantify the compensation performance, for each drift electric field from 0.1–2 kV/cm (with 0.1 kV/cm increments), $10^{4}$ $\nu_{e}$-Ar CC events are simulated in a LArTPC with a uniform neutrino energy distribution between 0.5–5 GeV. The $e/h$ is calculated for each E-field and plotted in the bottom panel of Fig. 7. The ideal compensation of $e/h$ close to 1 can be achieved from 0.3–0.5 kV/cm. The compensation is slightly worse at lower and higher E-fields, but good compensation of $e/h$ from 1–1.05 can still be achieved across a broad range of E-fields between 0.2 and 1.8 kV/cm.
@@
Figure 8: Simulated light calorimeter response of the hadronic component $h$ in LArTPC at the drift E-field of 0.5 kV/cm for different incident neutrino energies.
@@
The compensation performance of the LArTPC light calorimetry is also energy dependent. To study the energy dependence, the events in Fig. 5 are divided into different incident neutrino energy bins, and the calorimeter response of the hadronic component ($R_{\mathrm{cal}}^{h}$) in the light calorimetry is plotted in Fig. 8. For neutrino energies above 2 GeV, the $R_{\mathrm{cal}}^{h}$ distributions are similar, with smaller fluctuations for higher energy neutrinos. The $R_{\mathrm{cal}}^{h}$ distribution, however, is much wider for the 1 GeV or lower-energy incident neutrinos, which leads to a worse reconstructed energy resolution for lower-energy neutrinos (cf. Fig. 10) using light calorimetry.
@@
Figure 9: Simulated light calorimeter response ($R_{\mathrm{cal}}$) for protons, neutrons, and charged pions in LArTPC at the drift E-field of 0.5 kV/cm . The top panel shows a scatter plot of the available energy of each particle versus its $R_{\mathrm{cal}}$. The bottom panel shows the projected distribution of $R_{\mathrm{cal}}$ of different particles for all energies.
@@
This energy dependence is further studied in Fig. 9 by separately plotting the $R_{\mathrm{cal}}$ for protons, neutrons, and charged pions in the light calorimetry. The top panel shows a scatter plot of the available energy of each particle (i.e. the kinetic energy of protons and neutrons, and the total energy of the charged pions) versus its $R_{\mathrm{cal}}$. The bottom panel shows the projected distribution of $R_{\mathrm{cal}}$ of different particles for all energies. We can observe that at low energy below $\sim$1 GeV, the behavior is quite different between $p$, $n$, and $\pi^{\pm}$. This is because a large contribution of $R_{\mathrm{cal}}$ comes from the primary particle itself before it possibly starts a hadronic shower. The proton has a higher $R_{\mathrm{cal}}$ because it starts as a charged particle without missing energy, and receives extra recombination light. The neutron has a lower $R_{\mathrm{cal}}$ because it has a high missing energy from the beginning. The charged pion has a similar $dE/dx$ to a MIP particle before it starts the hardronic shower, so its $R_{\mathrm{cal}}$ is similar to electrons, with additional missing energy from particle decays. Therefore, the total $R_{\mathrm{cal}}$ of the hadronic component becomes dependent on the distribution of final state particles, leading to a wider distribution for lower energy neutrinos as seen in Fig. 8. However, at higher energy above $\sim$2 GeV, the $p$, $n$, and $\pi^{\pm}$ behave very similarly because they all start hadronic cascades and the $R_{\mathrm{cal}}$ is dominated by the secondary shower particles, which are similar for different primary particles. Thus, the compensation performance is better for higher-energy neutrinos and less sensitive to the modeling of final-state particles after neutrino interactions.
V Light calorimetry performance
@@
Based on the discussions in Sec. IV and Appendix B, in this section, we describe four energy reconstruction methods in LArTPC, one for light calorimetry and three for charge calorimetry, and compare their performance. The light calorimetry utilizes the self-compensating calorimeter concept described earlier. The charge calorimetry methods are similar in concept to previous works from LArTPC experiments and are included here for comparison.
@@
    L1: Simple light calorimetry by summing up $L$ from all particle activities and scaling to the incident neutrino energy,
@@
    Q1: Simple charge calorimetry by summing up $Q$ from all particle activities and scaling to the incident neutrino energy,
@@
    Q2: Assuming $e$ and $h$ charge activities can be grouped using pattern recognition algorithms on the 3D image (cf. Appendix B), scaling the $Q$ from $e$ and $h$ separately before summing up to the incident neutrino energy,
@@
We also studied two additional reconstruction methods but they are not reported in detail here. The first method is the $Q+L$ calorimetry based on Eq. 8. While it removes the recombination effect and recovers $E_{\mathrm{dep}}$, it does not account for the missing energy contribution, therefore, its performance is only slightly better than Q1 but worse than Q2 and L1. The second method assumes activities from individual particles can be grouped and their energies can be reconstructed separately. This is mostly useful for muons and charged pions since their missing energies from decays or captures can be corrected event by event, using methods such as a range-based energy reconstruction. The missing energy fluctuation from nuclear breakup still remains, e.g. from neutrons, making its performance only slightly better than Q3.
V.1 Energy resolution comparision
@@
Figure 10: (Top) Distribution of the reconstructed neutrino energy from $10^{4}$ simulated 3 GeV $\nu_{e}$-Ar CC interaction events in LArTPC at the drift E-field of 0.5 kV/cm. The four colored histograms correspond to the four reconstruction methods Q1–3 and L1, respectively. (Bottom) Relative energy resolution, $\sigma/\bar{E}$, as a function of the incident neutrino energy for the four reconstruction methods Q1–3 and L1.
@@
The top panel of Fig. 10 shows the distribution of the reconstructed neutrino energy from $10^{4}$ simulated 3 GeV $\nu_{e}$-Ar CC interactions using the Q1–3 and L1 reconstruction methods. Q1 (black) has a long tail extending to low energies, which comes from the missing energy contribution and the different calorimeter response between $e$ and $h$. L1 (red) removes the low-energy tail through the self-compensation mechanism, resulting in a more symmetric shape and much better resolution. Q2 (green) has a similar performance to L1, sharper near the peak but with a longer tail in the high energy. Q3 (blue) has the best overall performance. Both Q2 and Q3 utilize LArTPC’s charge imaging capability to separate $e$ and $h$ activities to improve the reconstructed energy resolution.
@@
The bottom panel of Fig. 10 shows the relative energy resolution, $\sigma/\bar{E}$, as a function of the incident neutrino energy. For each fixed neutrino energy, the $\sigma$ is calculated from the root mean square (RMS) of the distribution from each reconstruction method, and the mean energy of the distribution is used as $\bar{E}$. The RMS is larger than the width of the peak for a tailed distribution such as Q1 and Q2. This plot shows that Q3 achieves the best energy resolution at $\sim$6at 1 GeV and Q1 has the worst resolution of $\sim$14at 1 GeV. The L1 method has an 8.4% resolution at 1 GeV. The Q2 method performs similarly to L1, although the longer tail at high energy makes its RMS worse than L1. In practice, the 2D energy response matrices of reconstructed neutrino energy vs. true neutrino energy, also called the detector smearing matrices, are used in physics analysis, and they are shown in Fig. 14 in Appendix C. The energy resolution and response matrices from this study can be readily compared with other literature or experimental works [47, 40, 22, 48, 49], and can be used in the phenomenology studies such as those in Ref. [47, 50]. One should keep in mind the experimental conditions assumed in different studies.
@@
It’s worth noting that although Q3 has the best overall performance, the light calorimetry L1 is a simpler reconstruction method without needing complex pattern recognition algorithms to separate $e$ versus $h$, or track-like versus dot-like activities. For higher energy neutrinos above 3 GeV, the performance between L1 and the much more sophisticated Q3 methods is nearly equal owing to the self-compensation mechanism in the light calorimetry. Therefore, the light calorimetry in LArTPC provides an independent and comparable energy estimation, complementing the well-established charge calorimetry. In addition, since the light calorimetry through the self-compensation mechanism has very different systematic uncertainties compared to the charge-based calorimetry, such as its dependence on the final state particle distribution (cf. Fig. 9) from neutrino event generator models [42], it provides a new way of estimating and mitigating the systematic uncertainty in GeV neutrino energy reconstruction, which is one of the largest uncertainties affecting the measurements of neutrino oscillation parameters and $\nu$-Ar cross sections [51].
V.2 Impact of nonuniform light collection
In the previous studies, the light yield is assumed to be uniform across the entire detector volume, which in practice is difficult to achieve because of the constraints of photon detector placement and achievable surface coverage. Variation of light collection efficiency in large detectors presents a well-known challenge in scintillation-based energy reconstructions. Here we study the impact of nonuniform light collection on the energy resolution in LArTPC light calorimetry.
The nonuniformity of the photon collection efficiency (PCE) depends on specific detector designs. As an example, Fig. 15 of Ref. [24] shows a GEANT4 simulation of the APEX design for DUNE Phase-II far detectors. The average PCE in the APEX detector is 0.8%, leading to an average light yield of $\sim$180 photoelectrons per MeV. The photon detector locations are shown in Fig. 12 of Ref. [24]. While the light yield is rather uniform across the majority of the detector volume, it is smallest in the middle near the cathode and on the top and bottom near the anodes ($\sim$109 PE/MeV), and largest near the photon detectors at the field cage walls of the LArTPC ($\sim$300 PE/MeV). Given this variation of light collection, a simple light calorimetry reconstruction as described by Eq. 14 would not yield good results without position-dependent event selections, and additional steps are necessary to correct for the light collection nonuniformity. We note that efforts are ongoing to optimize the design of the APEX detector including methods such as adding photon detectors at the cathode and increasing the reflectivity of the anodes, all of which would further improve the uniformity of the light collection.
In an ideal event reconstruction, given a known photon collection efficiency map $\epsilon(x,y,z)$, the light nonuniformity can be corrected by calculating a correction factor $C$ for each event to obtain the effective PCE of this event ($\epsilon_{\mathrm{eff}}=C\cdot\bar{\epsilon}$),
$C=\frac{\sum_{i}(dL)_{i}\epsilon_{i}}{\bar{\epsilon}L},$ (18)
where $i$ goes through each voxel position of the reconstructed energy deposition. In the numerator, $(dL)_{i}$ is the reconstructed visible light in voxel $i$ and $\epsilon_{i}$ is the PCE in voxel $i$. In the denominator, $L$ is the total visible light and $\bar{\epsilon}$ is the average PCE of all voxels. While $\epsilon_{i}$ can be obtained from detector simulation and a well-designed calibration program, the reconstruction of $(dL)_{i}$ is known to be difficult unless for simple event topologies such as point-like events (low-energy electrons), or line-like events (high-energy muons). For more complex event topologies such as the neutrino interaction shown in Fig. 13, the reconstruction of $(dL)_{i}$ at a fine position resolution is challenging using light signals alone.
One of the main advantages of the LArTPC is its superb 3D imaging capability with mm-scale position resolution using ionization charge signals. This charge-based event reconstruction provides a new way to perform light nonuniformity correction, which is not available in other scintillator detectors that only record light signals. For each event, we can calculate a different correction factor $C^{\prime}$ using charge-based reconstruction:
$C^{\prime}=\frac{\sum_{i}(dQ)_{i}\epsilon_{i}}{\bar{\epsilon}Q},$ (19)
where $(dL)_{i}$ is replaced by $(dQ)_{i}$ — the reconstructed visible charge in each voxel $i$. Correspondingly, the total visible light $L$ is replaced by the total visible charge $Q$. Since $(dQ)_{i}$ can be well reconstructed with high position resolution [52, 53, 54, 55], $C^{\prime}$ can be calculated to a much better precision than $C$. If $(dQ)_{i}/(dL)_{i}$ is a constant, then $C^{\prime}$ would equal $C$ and the light nonuniformity can be perfectly corrected. However, since $(dQ)_{i}$ and $(dL)_{i}$ have a different dependence on $dE/dx$ (cf. Fig. 2), there is residual nonunformity after using Eq. 19, which needs to be quantified.
Figure 11: (Top) Reconstructed neutrino energy distributions using light calorimetry from $10^{4}$ simulated 3 GeV $\nu_{e}$-Ar CC interactions with 50% variation of position-dependent photon collection efficiency. The blue histogram assumes an average PCE following Eq. 14 without any nonuniformity corrections. The black histogram uses an ideal nonuniformity correction following Eq. 18. The red histogram uses the charge-based nonuniformity correction following Eq. 19. (Bottom) Reconstructed neutrino energy resolution of 3 GeV $\nu_{e}$-Ar CC interactions from light calorimetry using charge-based nonuniformity correction versus the size of the nonuniformity.
[…]